# Unsupervised Meta-Learning via In-Context Learning

### v2

@@ A.1 Experimental details
We compare our results with UML methods, an unsupervised few-shot learning method, a supervised meta-learning method, and two SSL methods. For the UML baselines, we consider CACTUs-MAML [19], CACTUs-ProtoNet [19], UMTRA [21], Meta-GMVAE [26], and PsCo [20]. All these methods are evaluated in-domain, i.e., using the same dataset for training and evaluation, to adhere to the setting proposed in the original papers. Only PsCo is also extended to the cross-domain scenario that we discuss in this paper. All methods are trained for $100$ epochs, using the parameters reported in the original papers [21, 19, 26, 20], and evaluated with $100$ adaptation steps on each task when required by the model. When evaluated in-domain, all approaches use a Conv5 architecture consisting of $5$ convolutional layers with $64$ filters and a kernel size of $3$, followed by batch normalization, ReLU non-linearity, max pooling, and a classifier head. The only exception is Meta-GMVAE. For this method, the authors trained a Conv5 feature extractor with SimCLR and input the learned features into a variational autoencoder (VAE) [26]. Due to time limitations and the computational resources required to train a model with SimCLR, in our experiments, we used a feature extractor consisting of a pre-trained version of SimCLR on ResNet-50 [3] using ImageNet-1k, followed by a projection layer fine-tuned for $100$ steps to the training dataset, as done for the SSL baselines. This approach results in a better performance than the one reported in the original paper [26] (see Table 1), likely due to the improved ability of the feature extractor to extract meaningful features. For PsCo, when evaluated on a cross-domain setting, we utilized the ResNet-50 architecture trained on ImageNet-964 to avoid data leakage, and we then applied the model to the test domain using $100$ adaptation steps to it. We also included BECLR [34] in our comparison utilizing the same hyperparameters and model architectures proposed in the original paper, given the importance of hyperparameter choice for final performance. Specifically, the ResNet-50 feature extractor was trained only on miniImageNet and evaluated on cross-domain scenarios in Table 2.
To compare our results with CAML [12], the same architecture and hyperparameter of our approach are applied to this method. This results in a lower performance for CAML compared to the original paper [12], as the reduced size of the model and the different feature extractor (ResNet-50 instead of ViT-CLIP), but it guarantees fair comparisons and lets us train the model with the available computational resources.
We also provide a comparison with two SSL approaches - SimCLR [7] and SwAV [6]. The details for training and evaluation are provided in Section 4.6.
@@ A.2 In-context learning analysis
Figure 4: Visualization of clustered embeddings obtained with CAMeLU after the fixed feature extractor (left) and the transformer encoder (right) across different datasets. The plots represent 5-way 5-shot tasks during inference. Crosses indicate the centroids for each class, triangles represents the query sample embeddings, and the numbers denote the Euclidean distances between the query and each class centroid. The plots are obtained using t-SNE [46] with a perplexity equals to 9.
To verify the contribution of the in-context learner in CAMeLU, we examine the embedding space learned during inference, both after the feature extractor and the transformer encoder. Figure 4 presents a t-SNE visualization of a single test task, where clusters represent the embeddings of the support classes. For simplicity, we illustrate a 5-way 5-shot task with one query sample for each dataset and we report the Euclidean distance between the query and the centroid of each class. As the dataset complexity increases, we observe greater variation between the embeddings learned from the fixed feature extractor and those refined by the transformer encoder. For instance, with the miniImageNet dataset, the feature extractor alone is able to recognize the class of the query, clustering the query sample with the support samples belonging to the same class. However, in more challenging datasets such as CUB and Aircraft, the embedding space after the feature extractor appears more sparse, reflected by the large Euclidean distance between the query sample and the centroid of each class. In contrast, the transformer encoder significantly improves the representation, producing more compact and well-separated clusters, underscoring its crucial role in CAMeLU. For instance, in the Aircraft dataset, the query would be misclassified as class 5 (in grey) based on the feature extractor alone, but it is correctly classified after passing through the transformer. This highlights the role of the transformer encoder in updating the support and query representations based on the context of the task, not only the image context, improving classification accuracy.
@@ A.3 Multi-dataset training
Table 5: Comparison of CAMeLU and CAML when trained on the single ImageNet-964 dataset (IN-964) and on a multi-dataset consisting of ImageNet-964 + MSCOCO + Fungi (mds). Results show the mean and standard deviations for the 5-way 1-shot and the 5-way 5-shot settings across three complete runs of the algorithms.
@@
We conduct additional experiments to evaluate the performance of CAMeLU when trained on a large-scale dataset. As our focus is on cross-domain classification through in-context learning, we hypothesize that training on a dataset spanning various concepts could enhance classification performance, as suggested in [12, 32]. To test this hypothesis, we combine three training datasets with varying levels of granularity: ImageNet-964 [8], MSCOCO [28], and Fungi [40]. During each training episode, a dataset is uniformly sampled, and $N$ data points are extracted. These samples are then utilized in our method as described in Section 3. Table 5 presents the results for CAMeLU and the supervised CAML method. Notably, training with a combination of multiple datasets (denoted as mds in Table 5) yields improved performance for CAML (supervised) compared to training solely on ImageNet-964, likely due to the increased variability in the sampled tasks. However, CAMeLU does not exhibit a similar performance boost, as its task creation mechanism already introduces substantial variability, reducing the benefit of additional dataset diversity. Moreover, the large size of ImageNet-964 (around $80\%$ of the overall dataset), leads to its more frequent selection compared to the other datasets and limits the potential for performance gains from the other datasets. Consequently, to optimize computational resources and time, we conducted all the other experiments by training solely on ImageNet-964.
@@ A.4 Ablation studies - Feature extractor
To assess the impact of the pre-trained feature extractor, we evaluate CAMeLU with various extractors pre-trained using different strategies. In particular, we compare the effectiveness of a ResNet-50 encoder pre-trained in a supervised manner both on ImageNet-1k and on ImageNet-964, a ResNet-50 pre-trained on ImageNet-1k using two SSL strategies, i.e., SimCLR [7] and SwAV [6], and a ViT-B/16 architecture pre-trained with CLIP [36] on a large dataset with $400$ million (image, text) pairs. The extractors were downloaded from the Hugging Face [57] and PyTorch Lightning Bolts [3] websites, except for the ResNet-50 architecture pre-trained on ImageNet-964. Results in Table 6(a) demonstrate that the models pre-trained on ImageNet-1k exhibit significantly higher performance on miniImageNet, compared to other datasets, primarily due to the data leakage issue described in Section 4.2. This issue was mitigated by pre-training the ResNet-50 architecture on ImageNet-964, which results in a drop in the performance on miniImageNet and CIFAR-fs due to the removal of data leakage, making it comparable with the results on CLIP-ViT-B/16. Notably, CLIP-ViT-B/16 stands out as the best performing method, due to its dataset-agnostic nature and its ability to learn representations that generalize across a broad range of tasks. Furthermore, when training CAMeLU with a multi-dataset approach (Table 6(b)), as described in Appendix A.3, results for CLIP-ViT-B/16 improve further, highlighting its applicability also to datasets significantly different from those used for training [36]. These findings demonstrate that CAMeLU’s performance scales with the strength of the feature extractor, indicating potential for further investigation as more robust feature extractors become available. However, in this work, we decided to use the ResNet-50 architecture to obtain fair comparison with previous baselines.
@@ A.5 Evaluation of the task creation mechanism
Figure 5: Validation accuracy on miniImageNet (mini) and CUB while training CAMeLU with two different task creation mechanisms. Red and purple curves are obtained with our proposed strategy in Section 3.1, while orange and pink curves are obtained by applying only data augmentations based on image manipulations to generate the support and query samples. The training is performed using ImageNet-964.
@@
To assess the effectiveness of the task creation strategy employed in our proposed approach, we conduct a comparative analysis of CAMeLU’s performance under two different task creation mechanisms. Specifically, we evaluate CAMeLU when tasks are generated solely using data augmentations for both the support and query sets, following a strategy similar to UMTRA [21], versus employing our proposed approach outlined in Section 3.1. The results presented in Table 7(a) across all the datasets suggest that our proposed strategy enhances the generalization on cross-domain datasets such as CUB, Aircraft, and Meta-iNat. This conclusion is further supported by Figure 5, which show the validation accuracy on miniImageNet and CUB using the two mechanisms discussed before, along with the CLIP-ViT feature extractor described in Appendix A.4. These results confirm that our task creation strategy is more robust, particularly in cross-domain evaluations, even when stronger feature extractors are utilized.
@@
Additionally, we experimented with k-means clustering as an alternative to the proposed task creation strategy. Inspired by CACTUs [19], we applied clustering on the embeddings generated by the feature extractor to generate pseudo-labels. The results are presented in Table 7(b) when training on miniImageNet due to the high computational cost of k-means and indicate that our proposed mechanism generalizes better across domains. Furthermore, k-means clustering requires an insight on the number of classes in the training dataset, as choosing a high number of clusters would lead to a lack of samples per class, whereas a low number may hinder generalization. In contrast, CAMeLU does not rely on such assumptions, enhancing its robustness compared to clustering-based approaches.
@@ A.6 Query samples generation strategy
We also conducted additional experiments to validate the choice of utilizing Equation 1 for generating query samples. In particular, we compare the results obtained as a linear combination of the augmented image $\tilde{x}_{n,j}$ with the randomly sampled image $z_{j}$ (pixel level), as in Equation 1, and at the patch level, as in [59]. Specifically, for the latter, we randomly select a patch from $z_{j}$ with an area ratio proportional to $\lambda$, and we paste it into $\tilde{x}_{n,j}$. Figure 6 illustrates an example of these two techniques by mixing two images sampled from ImageNet-964. As shown in Figure 6(c), merging the images at the pixel level results in a mixed image where some information from $\tilde{x}_{n,j}$ and $z_{j}$ is retained in every part of the image. Contrarily, in Figure 6(d), there is no information about $\tilde{x}_{n,j}$ in the lower left corner, forcing the network to attend only to the upper right part of the image to classify it with the same class as $\tilde{x}_{n,j}$. Therefore, we hypothesize that the pixel level strategy is more suitable for our approach as the goal is to attend to the whole image to extract robust features that allow the model to classify the query image with the same class as the support one while ensuring diversity between the two. To validate this, we utilized the Structural Similarity Index ($SSIM$) [53]. $SSIM$ is used as a metric to measure the similarity between two given images based on three image features: luminance, contrast, and structure. Formally, considering $x$ and $y$ two given images, $SSIM$ is calculated as follows:
@@
Instead of applying the above formula all over the image at once, [53] proposed a local variant that consists of computing the $SSIM$ index locally and averaging these values to obtain the global $SSIM$ value. For our purpose, we computed this metric between each of the two images used for the generation, i.e., $\tilde{x}_{n,j}$ and $z_{j}$, and the resulting mixed image, i.e., $x_{j}^{(qr)}$. We then average the results, obtaining an indicator, denoted as $mSSIM$, of how similar the query image is with the images used for the generation, or in other words, how much local information is retained from $\tilde{x}_{n,j}$ and $z_{j}$ into $x_{j}^{(qr)}$. Results are shown in Table 8 for $\lambda=0.25$ and $\lambda=0.49$, confirming the hypothesis that, even for high $\lambda$ values, mixing at the pixel level guarantees more information retained across the whole image compared to using a patch level strategy. This is also confirmed by the results in Table 9, which shows a decrease in the performance when CAMeLU is trained with the patch level strategy for query generation.
@@
We also ablate the values of the $\alpha$ and $\beta$ parameters used in the $Beta$ distribution from which $\lambda$ is sampled. Table 9 presents the results for different values of $\alpha$ and $\beta$ when $\lambda\sim Beta(\alpha,\beta)$ and $\lambda\in(0,0.5)$. The results indicate that the optimal choice for CAMeLU is to select $\alpha=1,\beta=1$, which appears to be a uniform distribution. Additionally, $\alpha=2,\beta=5$ also yields comparable results, highlighting the importance of selecting a sufficiently small $\lambda$ to ensure the incorporation of sufficient information from $\tilde{x}_{n,j}$ into $x_{j}^{(qr)}$ facilitating the model’s ability to classify the latter with the same class as $x_{n,i}^{(sp)}$.
Table 9: Accuracy results of CAMeLU when trained with different strategies for generating the query samples. Pixel level mix refers to the scenario where query samples are generated with Equation 1, while patch level mix refers to a strategy similar to the one proposed in [59]. Results are reported in the 5-way 5-shot scenario with $\lambda\sim Beta(\alpha,\beta)$ and different $\alpha$ and $\beta$ values. Results show the mean and standard deviations across three complete runs of the algorithms.
@@
A.7 Analysis of computational complexity
The core of CAMeLU relies on a transformer architecture, which is known for its computational demands due to the self-attention mechanism. The transformer model has a computational complexity of $O(n^{2}\cdot d)$ per layer [48], where $n$ is the sequence length and $d$ is the hidden dimension. In our context, $n$ includes both the support samples and the query sample. Consequently, the total computational complexity for evaluating $Q$ queries is $O(Q\cdot(NK+1)^{2}\cdot d)$, where $N$ is the number of classes, $K$ the number of shots, $NK+1$ indicates one query per input sequence, and $Q$ is the total number of queries. This results in a quadratic complexity in the number of support samples which can be computationally demanding. However, we have demonstrated in Table 1 that CAMeLU achieves good performance even with only $K=1$ support sample per class. Additionally, further experiments with only one query sample per episode and one support sample per class (i.e., $N=5$, $K=1$, $Q=1$) yield results of $80.74\pm 0.65$ for miniImageNet, $63.07\pm 1.14$ for CIFAR-fs, $54.84\pm 1.41$ for CUB, $30.32\pm 0.76$ for Aircraft, and $55.76\pm 0.08$ for Meta-iNat. These results are comparable to those reported in Table 1 for the 5-ways 1-shot scenario using 25 queries, highlighting that a single query is sufficient for good performance, as also demonstrated in CAML.

### v3

@@ A.1 Experimental details
We compare our results with UML methods, an unsupervised few-shot learning method, a supervised meta-learning method, and two SSL methods. For the UML baselines, we consider CACTUs-MAML (Hsu et al., 2018), CACTUs-ProtoNet (Hsu et al., 2018), UMTRA (Khodadadeh et al., 2019), Meta-GMVAE (Lee et al., 2020), and PsCo (Jang et al., 2022). All these methods are evaluated in-domain, i.e., using the same dataset for training and evaluation, to adhere to the setting proposed in the original papers. Only PsCo is also extended to the cross-domain scenario that we discuss in this paper. All methods are trained for $100$ epochs, using the parameters reported in the original papers (Khodadadeh et al., 2019; Hsu et al., 2018; Lee et al., 2020; Jang et al., 2022), and evaluated with $100$ adaptation steps on each task when required by the model. When evaluated in-domain, all approaches use a Conv5 architecture consisting of $5$ convolutional layers with $64$ filters and a kernel size of $3$, followed by batch normalization, ReLU non-linearity, max pooling, and a classifier head. The only exception is Meta-GMVAE. For this method, the authors trained a Conv5 feature extractor with SimCLR and input the learned features into a variational autoencoder (VAE) (Lee et al., 2020). Due to time limitations and the computational resources required to train a model with SimCLR, in our experiments, we used a feature extractor consisting of a pre-trained version of SimCLR on ResNet-50 (Borovec et al., 2022) using ImageNet-1k, followed by a projection layer fine-tuned for $100$ steps to the training dataset, as done for the SSL baselines. This approach results in a better performance than the one reported in the original paper (Lee et al., 2020) (see Tab. 1), likely due to the improved ability of the feature extractor to extract meaningful features. For PsCo, when evaluated on a cross-domain setting, we utilized the ResNet-50 architecture trained on ImageNet-964 to avoid data leakage, and we then applied the model to the test domain using $100$ adaptation steps to it. We also included BECLR (Poulakakis-Daktylidis & Jamali-Rad, 2024) in our comparison utilizing the same hyperparameters and model architectures proposed in the original paper, given the importance of hyperparameter choice for final performance. Specifically, the ResNet-50 feature extractor was trained only on miniImageNet and evaluated on cross-domain scenarios in Tab. 2.
To compare our results with CAML (Fifty et al., 2024), the same architecture and hyperparameter of our approach are applied to this method. This results in a lower performance for CAML compared to the original paper (Fifty et al., 2024), as the reduced size of the model and the different feature extractor (ResNet-50 instead of ViT-CLIP), but it guarantees fair comparisons and lets us train the model with the available computational resources.
We also provide a comparison with two SSL approaches - SimCLR (Chen et al., 2020) and SwAV (Caron et al., 2020). The details for training and evaluation are provided in Sect. 4.6.
@@ A.2 In-context learning analysis
Figure 5: Visualization of clustered embeddings obtained with CAMeLU after the fixed feature extractor (left) and the transformer encoder (right) across different datasets. The plots represent 5-way 5-shot tasks during inference. Crosses indicate the centroids for each class, triangles represent the query sample embeddings, and the numbers denote the Euclidean distances between the query and each class centroid. The plots are obtained using t-SNE (Van der Maaten & Hinton, 2008) with a perplexity equals to 9.
To verify the contribution of the in-context learner in CAMeLU, we examine the embedding space learned during inference, both after the feature extractor and the transformer encoder. Fig. 5 presents a t-SNE visualization of a single test task, where clusters represent the embeddings of the support classes. For simplicity, we illustrate a 5-way 5-shot task with one query sample for each dataset, and we report the Euclidean distance between the query and the centroid of each class. As the dataset complexity increases, we observe greater variation between the embeddings learned from the fixed feature extractor and those refined by the transformer encoder. For instance, with the miniImageNet dataset, the feature extractor alone is able to recognize the class of the query, clustering the query sample with the support samples belonging to the same class. However, in more challenging datasets such as CUB and Aircraft, the embedding space after the feature extractor appears more sparse, reflected by the large Euclidean distance between the query sample and the centroid of each class. In contrast, the transformer encoder significantly improves the representation, producing more compact and well-separated clusters, underscoring its crucial role in CAMeLU. For instance, in the Aircraft dataset, the query would be misclassified as class 5 (in grey) based on the feature extractor alone, but it is correctly classified after passing through the transformer. This highlights the role of the transformer encoder in updating the support and query representations based on the context of the task, not only the image context, improving classification accuracy.
@@ A.3 Multi-dataset training
Table 5: Comparison of CAMeLU and CAML when trained on the single ImageNet-964 dataset (IN-964) and on a multi-dataset (mds) consisting of ImageNet-964 + MSCOCO + Fungi. Results show the mean and standard deviations for the 5-way 1-shot and the 5-way 5-shot settings across three complete runs of the algorithms.
@@
We conduct additional experiments to evaluate the performance of CAMeLU when trained on a large-scale dataset. As our focus is on cross-domain classification through in-context learning, we hypothesize that training on a dataset spanning various concepts could enhance classification performance, as suggested in Min et al. (2022) and Fifty et al. (2024). To test this hypothesis, we combine three training datasets with varying levels of granularity: ImageNet-964 (Deng et al., 2009), MSCOCO (Lin et al., 2014), and Fungi (Schroeder & Cui, 2018). During each training episode, a dataset is uniformly sampled, and $N$ data points are extracted. These samples are then utilized in our method as described in Sect. 3. Table 5 presents the results for CAMeLU and the supervised CAML method. Notably, training with a combination of multiple datasets (denoted as mds in Tab. 5) yields improved performance for CAML (supervised) compared to training solely on ImageNet-964, likely due to the increased variability in the sampled tasks. However, CAMeLU does not exhibit a similar performance boost, as its task creation mechanism already introduces substantial variability, reducing the benefit of additional dataset diversity. Moreover, the large size of ImageNet-964 (around $80\%$ of the overall dataset), leads to its more frequent selection compared to the other datasets and limits the potential for performance gains from the other datasets. Consequently, to optimize computational resources and time, we conducted all the other experiments by training solely on ImageNet-964.
@@ A.4 Ablation studies - Feature extractor
To assess the impact of the pre-trained feature extractor, we evaluate CAMeLU with various extractors pre-trained using different strategies. In particular, we compare the effectiveness of a ResNet-50 encoder pre-trained in a supervised manner both on ImageNet-1k and on ImageNet-964, a ResNet-50 pre-trained on ImageNet-1k using two SSL strategies, i.e., SimCLR (Chen et al., 2020) and SwAV (Caron et al., 2020), and a ViT-B/16 architecture pre-trained with CLIP (Radford et al., 2021) on a large dataset with $400$ million (image, text) pairs. The extractors were downloaded from the Hugging Face (Wolf et al., 2019) and PyTorch Lightning Bolts (Borovec et al., 2022) websites, except for the ResNet-50 architecture pre-trained on ImageNet-964. Results in Tab. 6(a) demonstrate that the models pre-trained on ImageNet-1k exhibit significantly higher performance on miniImageNet, compared to other datasets, primarily due to the data leakage issue described in Sect. 4.2. This issue is mitigated by pre-training the ResNet-50 architecture on ImageNet-964, which results in a drop in the performance on miniImageNet and CIFAR-fs due to the removal of data leakage, making it comparable with the results on CLIP-ViT-B/16. For CLIP-ViT-B/16, however, thorough verification of potential data leakage was not possible due to the undisclosed nature of its training dataset. As such, these results should be interpreted with caution. CLIP-ViT-B/16 stands out as the best-performing method due to its dataset-agnostic nature and its ability to learn representations that generalize across a broad range of tasks. Furthermore, when training CAMeLU with a multi-dataset approach (Tab. 6(b)), as described in Appendix A.3, results for CLIP-ViT-B/16 improve further, highlighting its applicability also to datasets significantly different from those used for training (Radford et al., 2021). These findings demonstrate that CAMeLU’s performance scales with the strength of the feature extractor, indicating potential for further investigation as more robust feature extractors become available. However, in this work, we decided to use the ResNet-50 architecture to guarantee a fair comparison with previous baselines.
@@ A.5 Evaluation of the task creation mechanism
Figure 6: Validation accuracy on miniImageNet (mini) and CUB while training CAMeLU with two different task creation mechanisms. Red and purple curves are obtained with our proposed strategy in Sect. 3.1, while orange and pink curves are obtained by applying only data augmentations based on image manipulations to generate the support and query samples. The training is performed using ImageNet-964.
@@
To assess the effectiveness of the task creation strategy employed in our proposed approach, we conduct a comparative analysis of CAMeLU’s performance under two different task creation mechanisms. Specifically, we evaluate CAMeLU when tasks are generated solely using data augmentations for both the support and query sets, following a strategy similar to UMTRA (Khodadadeh et al., 2019), versus employing our proposed approach outlined in Sect. 3.1. By applying our task creation strategy, we generate more complex tasks, making the generalization problem harder and the in-context learner more robust (Chan et al., 2022; Singh et al., 2024). The results presented in Tab. 7(a) confirm our claim. Across all the datasets, our proposed strategy enhances the generalization on cross-domain datasets such as CUB, Aircraft, and Meta-iNat. This conclusion is further supported by Fig. 6, which shows the validation accuracy on miniImageNet and CUB using the two mechanisms discussed before, along with the CLIP-ViT feature extractor described in Appendix A.4. These results confirm that our task creation strategy is more robust, particularly in cross-domain evaluations, even when stronger feature extractors are utilized.
@@
Additionally, we experimented with k-means clustering as an alternative to the proposed task creation strategy. Inspired by CACTUs (Hsu et al., 2018), we applied clustering on the embeddings generated by the feature extractor to generate pseudo-labels. The results are presented in Tab. 7(b) when training on miniImageNet due to the high computational cost of k-means and indicate that our proposed mechanism generalizes better across domains. Furthermore, k-means clustering requires an insight into the number of classes in the training dataset, as choosing a high number of clusters would lead to a lack of samples per class, whereas a low number may hinder generalization. In contrast, CAMeLU does not rely on such assumptions, enhancing its robustness compared to clustering-based approaches.
Finally, we show that the benefits of our task creation strategy extend beyond CAMeLU. In Tab. 8 we apply our proposed mechanism to generate pseudo-tasks on top of MAML (Finn et al., 2017). This allows for a direct comparison with MAML-based approaches, such as UMTRA (Khodadadeh et al., 2019) and CACTUs-MAML (Hsu et al., 2018), by replacing their original task creation mechanisms. The increased performance of our strategy applied to the baselines demonstrates its superiority over previous task creation methods. It may be objected that CACTUs-MAML achieves higher performance on miniImageNet and CIFAR-fs. However, this is caused by the use of a feature extractor pre-trained on ImageNet-1k, which introduces an unfair advantage by leaking information about the test data into the training phase. This performance gap narrows when the test distribution deviates from the training data (e.g., CIFAR-fs) and disappears for datasets with low correlation to ImageNet-1k, supporting our hypothesis. Indeed, on datasets that share low similarity with ImageNet-1k, our method consistently outperforms both UMTRA and CACTUs-MAML. While these results highlight the strength of our task creation strategy, the performance still remains significantly lower than CAMeLU, especially in cross-domain scenarios. This emphasizes the critical role of combining our robust task creation mechanism with the in-context learning capabilities of transformer-based architectures to achieve superior performance.
@@
Table 8: Evaluation of the proposed task creation strategy when applied to build pseudo-tasks on top of MAML. The results are compared with other MAML-based baselines for UML, such as UMTRA and CACTUs-MAML.
miniImageNet CIFAR-fs CUB Aircraft Meta-iNat Method 5w1s 5w5s 5w1s 5w5s 5w1s 5w5s 5w1s 5w5s 5w1s 5w5s CACTUs-MAML $\mathbf{43.30}$ $\mathbf{54.21}$ $\mathbf{42.00}$ $\mathbf{56.64}$ $31.19$ $36.81$ $24.06$ $27.26$ $20.13$ $21.84$ UMTRA $39.93$ $50.73$ $32.93$ $46.13$ $27.06$ $36.60$ $22.40$ $31.73$ $28.96$ $37.12$ Proposed + MAML $34.04$ $46.13$ $37.02$ $52.00$ $\mathbf{31.22}$ $\mathbf{41.54}$ $\mathbf{26.12}$ $\mathbf{34.34}$ $\mathbf{30.94}$ $\mathbf{42.43}$
@@ A.6 Query samples generation strategy
We also conducted additional experiments to validate the choice of utilizing Eq. 1 for generating query samples. In particular, we compare the results obtained as a linear combination of the augmented image $\tilde{x}_{n,j}$ with the randomly sampled image $z_{j}$ (pixel level), as in Eq. 1, and at the patch level, as in Yun et al. (2019). Specifically, for the latter, we randomly select a patch from $z_{j}$ with an area ratio proportional to $\lambda$, and we paste it into $\tilde{x}_{n,j}$. Fig. 7 illustrates an example of these two techniques by mixing two images sampled from ImageNet-964. As shown in Fig. 7(c), merging the images at the pixel level results in a mixed image where some information from $\tilde{x}_{n,j}$ and $z_{j}$ is retained in every part of the image. Contrarily, in Fig. 7(d), there is no information about $\tilde{x}_{n,j}$ in the lower left corner, forcing the network to attend only to the upper right part of the image to classify it with the same class as $\tilde{x}_{n,j}$. Therefore, we hypothesize that the pixel level strategy is more suitable for our approach as the goal is to attend to the whole image to extract robust features that allow the model to classify the query image with the same class as the support one while ensuring diversity between the two. To validate this, we utilized the Structural Similarity Index ($SSIM$) (Wang et al., 2004). $SSIM$ is used as a metric to measure the similarity between two given images based on three image features: luminance, contrast, and structure. Formally, considering $x$ and $y$ two given images, $SSIM$ is calculated as follows:
@@
Instead of applying the above formula all over the image at once, (Wang et al., 2004) proposed a local variant that consists of computing the $SSIM$ index locally and averaging these values to obtain the global $SSIM$ value. For our purpose, we computed this metric between each of the two images used for the generation, i.e., $\tilde{x}_{n,j}$ and $z_{j}$, and the resulting mixed image, i.e., $x_{j}^{(qr)}$. We then average the results, obtaining an indicator, denoted as $mSSIM$, of how similar the query image is with the images used for the generation, or in other words, how much local information is retained from $\tilde{x}_{n,j}$ and $z_{j}$ into $x_{j}^{(qr)}$. Results are shown in Tab. 9 for $\lambda=0.25$ and $\lambda=0.49$, confirming the hypothesis that, even for high $\lambda$ values, mixing at the pixel level guarantees more information retained across the whole image compared to using a patch level strategy. This is also confirmed by the results in Tab. 10, which shows a decrease in the performance when CAMeLU is trained with the patch level strategy for query generation.
@@
We also ablate the values of the $\alpha$ and $\beta$ parameters used in the $Beta$ distribution from which $\lambda$ is sampled. Tab. 10 presents the results for different values of $\alpha$ and $\beta$ when $\lambda\sim Beta(\alpha,\beta)$ and $\lambda\in(0,0.5)$. The results indicate that the optimal choice for CAMeLU is to select $\alpha=1,\beta=1$, which appears to be a uniform distribution. Additionally, $\alpha=2,\beta=5$ also yields comparable results, highlighting the importance of selecting a sufficiently small $\lambda$ to ensure the incorporation of sufficient information from $\tilde{x}_{n,j}$ into $x_{j}^{(qr)}$ facilitating the model’s ability to classify the latter with the same class as $x_{n,i}^{(sp)}$.
Table 10: Accuracy results of CAMeLU when trained with different strategies for generating the query samples. Pixel level mix refers to the scenario where query samples are generated with Eq. 1, while patch level mix refers to a strategy similar to the one proposed in Yun et al. (2019). Results are reported in the 5-way 5-shot scenario with $\lambda\sim Beta(\alpha,\beta)$ and different $\alpha$ and $\beta$ values. Results show the mean and standard deviations across three complete runs of the algorithms.
@@
A.7 Computational complexity and resources usage
Table 11: Computational and time complexity of CAMeLU in comparison with PsCo. The comparison is performed considering the time required for the task creation, the training time (expressed in time per epoch), the inference time on a single task, and the GPU and CPU memory usage during training and inference.
Time task construction Training time Inference time
(ms) (ms/epoch) (ms/task)
PsCO $20772$ $4613656$ $605$
CAMeLU $1376$ $153000$ $57$
GPU training CPU training GPU inference CPU inference
(MiB) (MiB) (MiB) (MiB)
PsCO $43904$ $20904$ $1630$ $2061$
CAMeLU $6250$ $2588$ $3224$ $1667$
In this Sect., we analyze the computational and time complexity of CAMeLU and we compare it with PsCo Jang et al. (2022). Tab. 11 presents the time required for task generation, model training, and inference, along with GPU and CPU memory usage. The results demonstrate that CAMeLU is not only faster than PsCo but also significantly more memory efficient. Notably, CAMeLU requires only $57$ms for task inference, making it particularly suitable for real-time applications.
The computational complexity of CAMeLU is primarily attributed to the transformer architecture, which is known for its computational demands due to the self-attention mechanism. The transformer model has a computational complexity of $O(n^{2}\cdot d)$ per layer (Vaswani et al., 2017), where $n$ is the sequence length and $d$ is the hidden dimension. In our context, $n$ includes both the support samples and the query sample. Consequently, the total computational complexity for evaluating $Q$ queries is $O(Q\cdot(NK+1)^{2}\cdot d)$, where $N$ is the number of classes, $K$ the number of shots, $NK+1$ indicates one query per input sequence, and $Q$ is the total number of queries. This results in a quadratic complexity in the number of support samples which can be computationally demanding. However, we have demonstrated in Tab. 1 that CAMeLU achieves good performance even with only $K=1$ support sample per class. Additionally, further experiments with only one query sample per episode and one support sample per class (i.e., $N=5$, $K=1$, $Q=1$) yield results of $80.74\pm 0.65$ for miniImageNet, $63.07\pm 1.14$ for CIFAR-fs, $54.84\pm 1.41$ for CUB, $30.32\pm 0.76$ for Aircraft, and $55.76\pm 0.08$ for Meta-iNat. These results are comparable to those reported in Tab. 1 for the 5-ways 1-shot scenario using 25 queries, highlighting that a single query is sufficient for good performance, as also demonstrated in CAML.